OpenStudy (anonymous):

if 5/8 of a number is 3 less than 3/4 of the number,what is the number?

OpenStudy (angelab97):

make up the equation 5/8x= 3/4-3(x)

OpenStudy (anonymous):

mhm....

OpenStudy (angelab97):

5/8x+3x=3/4

OpenStudy (anonymous):

k thanx